Solve the equation xfrac{dy}{dx}-ay=x+1 where a is a constant
1
Expert's answer
2019-10-14T12:02:49-0400

First we find the general solution of "xy'-ay=0".

The chatacteristic equation of "xy'-ay=0" is "\lambda-a=0", which has solution "\lambda_0=a".

We obtain that the general solution of "xy'-ay=0" is "y=Cx^{\lambda_0}=Cx^a"

Next, we find the partial solution of "xy'-ay=x+1" of the form "y=Ax+B"

We have "xy'-ay=x(Ax+B)'-a(Ax+B)=A(1-a)x-aB".

So we obtain that if "a\neq 0" and "a\neq 1" , then "A(1-a)x-aB=x+1", that is "A(1-a)=1" and "-aB=1". Then the general solution of "xy'-ay=x+1" is "Cx^a+Ax+B=Cx^a+\frac{1}{1-a}x-\frac{1}{a}"

Consider the case "a=0". In this case we have the equation "xy'=x+1", that is "y'=1+\frac{1}{x}", so "y=\int\left(1+\frac{1}{x}\right)dx=x+\ln|x|+C".

Consider case "a=1". In this case we have the equation "x\frac{dy}{dx}-y=x+1", that is "xdy-ydx=(x+1)dx". Divide this equation by "x^2" and obtain "\frac{xdy-ydx}{x^2}=\left(\frac{1}{x}+\frac{1}{x^2}\right)dx"

"\frac{xdy-ydx}{x^2}=d\left(\frac{y}{x}\right)" and "\left(\frac{1}{x}+\frac{1}{x^2}\right)dx=d\left(\ln|x|-\frac{1}{x}\right)" , so we obtain the solution "\frac{y}{x}=\ln|x|-\frac{1}{x}+C", that is "y=x\ln|x|-1+Cx"

Answer:"y=\begin{cases}
Cx^a+\frac{1}{1-a}x-\frac{1}{a},&\text{if $a\not\in\{0,1\}$}\\
x+ln|x|+C,&\text{if $a=0$}\\
xln|x|-1+Cx,&\text{if $a=1$}
\end{cases}"